Grade 11Algebra

find the rth term of an A.P. .,SUM OF WHOSE FIRST n TERMS is 2n+3n^2

sharan
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
As we know
Tn = Sn - Sn-1
Similarily
Tr = Sr - Sr-1
Hence
Tr = 2r + 3r^2 - (2(r-1) + 3(r-1)^2)
Tr = 2r + 3r^2 -2r +2 - 3r^2 -3 + 6r
Tr = -1 +6r